# Towards Locally Deployable Fine-Tuned Causal Large Language Models for Mode Choice Behaviour

### v1

@@ 3.1 Instant-Level LLMs Mode Choice Predictive Performance Evaluation
Figure 3 presents the distribution of weighted F1 scores achieved by each model across three travel survey datasets; Swissmetro, Brightwater, and London PMC under three learning styles: zero-shot, random few-shot, and targeted few-shot. The violin plots show each model’s distribution across four experimental runs, highlighting central tendency (mean), maximum performance (peak), and dispersion (IQR). The vertical span of each violin reflects the full value range across runs. This structure allows us to simultaneously evaluate performance levels and consistency across models and prompting styles while being able to assess the model’s sensitivity to varying prompting styles and temperature. For the remainder of this section, the results are presented and discussed dataset-wise and summary tables for selected models, mean, peak and sensitivity/ consistency are listed accordingly.
@@
For the Swissmetro travel survey dataset, zero-shot experiments, DeepSeek R1 Distill Llama 8B outperforms others with a mean of 0.542 and peak of 0.568, maintaining a narrow IQR of 0.018, which indicates robust and insensitive performance under varying prompt conditions and temperature. Gemma 3 12B follows with a mean value of 0.497 and an even lower IQR of 0.011. Meanwhile, smaller models like Llama 3.2 1B record lower means (0.395) but also low dispersion (IQR = 0.011), suggesting output rigidity and predictably limited capacity at the base level, especially without any guiding examples.
@@
When transitioning to random few-shot learning, most models show gains in mean performance, though often at the cost of increased IQR as a result of the variability in the presented training profiles. DeepSeek R1 Distill Llama 8B reaches a new mean of 0.574 (peak = 0.617, IQR = 0.031), while Gemma 3 12B rises to 0.529 (peak = 0.558, IQR = 0.023), still maintaining less sensitivity to the variation of the presented training profiles, albeit higher than its baseline IQR performance.
@@
Targeted few-shot prompts significantly elevate both accuracy and stability. Gemma 3 12B reaches the highest overall mean of 0.619 (peak = 0.625), while tightening its IQR to just 0.0035, making it the most stable top performer. DeepSeek Llama follows with a mean = 0.601 but with a higher sensitivity to the prompt style and temperature. Table 2 summarizes the top model performance metrics for the Swissmetro dataset.
@@
Table 2: Swissmetro: Top model performance summary per learning style
@@
Zero‑Shot 0.542 (DeepSeek R1) 0.568 (DeepSeek R1) 0.011 (Gemma 3 12B)
Random Few‑Shot 0.574 (DeepSeek R1) 0.617 (DeepSeek R1) 0.023 (Gemma 3 12B)
Targeted Few‑Shot 0.619 (Gemma 3 12B) 0.625 (Gemma 3 12B) 0.0035 (Gemma 3 12B)
As for Brightwater, this travel survey dataset exhibits generally lower performance across all learning regimes. The complexity stems from its dynamic choice structure: while each individual sample presents only 4 alternatives, the global choice set comprises 10 distinct modes. This variability, combined with relatively homogeneous commuter profiles, renders the prediction task more challenging. In the zero-shot setting, Gemma 3 4B achieves the highest mean score (0.345) and peak performance (0.393), albeit with a comparatively wide IQR of 0.080, indicating moderate sensitivity to prompt and temperature variations. DeepSeek R1 Distilled models, on the other hand, deliver a lower mean (0.310) but exhibit greater consistency (Average IQR = 0.020),
With random few-shot learning, incremental gains are observed across the board. Gemma 3 12B attains the highest mean of 0.405 and peak of 0.412 while achieving a tight IQR of 0.010. Llama 3.2 1B shows smaller improvements (mean = 0.311, peak = 0.321) with consistent spread (IQR = 0.008 ).
Under the targeted few-shot, Llama 3 8B emerges as the top-performing model, achieving the highest mean (0.462) and peak (0.484) scores. However, in terms of stability, Gemma 3 4B proves superior, delivering a notably lower IQR of just 0.003 but with a lower mean of 0.384. Table 3 summarizes the top model performance metrics for the BrightWater dataset.
Table 3: Brightwater: Top model performance summary per learning style
Learning Style Top Mean (Model) Top Peak (Model) Tightest IQR (Model)
Zero‑Shot 0.345 (Gemma 3 4B) 0.393 (Gemma 3 4B) 0.008 (DeepSeek R1)
Random Few‑Shot 0.405 (Gemma 3 12B) 0.412 (Gemma 3 12B) 0.010 (Gemma 3 12B)
Targeted Few‑Shot 0.462 (Llama 3 8B) 0.484 (Llama 3 8B) 0.003 (Gemma 3 4B)
The analysis of model’s performance for London PMC shows relatively higher baseline performance. In the zero-shot setup, Mistral 7B leads with a mean of 0.576 and a peak of 0.584, with IQR = 0.006. Stealth 1.3 7B performs less consistently, showing a much broader IQR = 0.171 despite competitive peaks. Random few-shot prompts elevate Gemma 3 12B to a mean of 0.661 (peak = 0.675), with an IQR of only 0.007. Meanwhile, Stealth models appear competitive with relatively high peak values, although with much higher sensitivity to the examples and/ or the variation in the prompting styles and temperatures. In the targeted few-shot, Gemma 3 12B retains the highest mean (0.639) and a strong peak (0.647), while Qwen 2.5 7B hits the highest peak of 0.663 (mean = 0.622, IQR = 0.018). Stealth 1.3 7B remains remarkably consistent (IQR = 0.012), and respectable overall predictive performance showing great adaptability when prompted with tailored examples. Table 4 summarizes the top model performance metrics for the London PMC dataset.
Table 4: London PMC: Top model performance summary per learning style
Regime Top Mean (Model) Top Peak (Model) Tightest IQR (Model)
Zero‑Shot 0.576 (Mistral 7B) 0.584 (Mistral 7B) 0.006 (Mistral 7B)
Random Few‑Shot 0.661 (Gemma 3 12B) 0.675 (Gemma 3 12B) 0.007 (Gemma 3 12B)
Targeted Few‑Shot 0.639 (Gemma 3 12B) 0.663 (Qwen 2.5 7B) 0.012 (Stealth 1.3 7B)
@@
Across all three datasets, the transition from zero-shot to random few-shot prompting consistently yields notable improvements in mean weighted F1 scores, typically ranging from +0.03 to +0.07. However, this performance gain is frequently accompanied by an increase in interquartile range, with several models exhibiting expanded variability between runs. This pattern suggests heightened sensitivity to the quality and coherence of uncurated in-context examples, potentially undermining consistency despite improvements in central tendency. In contrast, targeted few-shot prompting not only enhances predictive accuracy but also systematically improves reliability. By selecting contextually relevant exemplars based on structured multi-layered similarity, IQR values are frequently reduced, often by 0.03 to 0.10, which highlights their stabilizing effect across models. This dual benefit affirms the critical role of example selection in few-shot learning and clearly demonstrates the importance of domain aware training examples.
l Refer to caption
@@
To further interrogate the variability observed in the weighted F1 scores and identify which experimental factors most strongly influence predictive performance, we decompose the weighted F1 variance across the four key dimensions: the choice of language model (Model), the learning strategy (Shot-Type: zero-shot vs. few-shot), the prompt style (Prompt-style), and the sampling Temperature (Temp).
Figure 4 presents the decomposition of the variance share that can be explained by each factor across the three travel survey datasets: Swissmetro, Brightwater, and London PMC. For each dataset, we fit an ordinary least squares (OLS) model of the form:
[…]
@@
We then conducted a Type II ANOVA to assess the relative contribution of each factor. The sum of squares attributable to each term was normalized such that the four contributions sum to 100% of the explained variance.
@@
In Swissmetro, the choice of model is the dominant driver of performance variability, accounting for 70.6% of the explained variance. This indicates that the choice of a more capable model yields the greatest returns. The Shot-Type contributes 29.0%, which constitutes a significant portion of the predictive power. Prompting style and temperatures impact compared to model choice and learning strategy are marginal, explaining just 0.36% and 0.10% respectively. The analysis of the Brightwater dataset, on the other hand, exhibits a reversed pattern, with Shot-Type explaining a substantial 79.8% of the variance. This reflects the increased difficulty of the task and the pivotal impact of in-context examples in elevating performance. Here, model choice plays a secondary role (19.9%), while Prompt-Type and Temperature again contribute minimally (0.21% and 0.12%). The findings suggest that for low-baseline tasks, few-shot examples can be transformative, whereas model upgrades alone yield a less effective boost in returns. As for the London PMC dataset, the analysis aligns more closely with Swissmetro. Model choice dominates (75.8% of variance), followed by Shot-Type (23.0%). Prompt-Type contributes slightly more here than in the other datasets (1.1%), though still marginal. Temperature influence is effectively null (0.005%). The results simply demonstrate the importance of selecting a performant LLM and a suitable learning strategy (zero- or few-shot), while suggesting that minor prompt adjustments or temperature tuning have limited practical value, and can be used for minor gains for achieving optimal performance.
@@
Overall, the analysis shows that for more traditional choice sets and diversified commuter profiles through larger samples, upgrading the model explains over 70% of F1 variation as we can note in both Swissmetro and London PMC. However, for more challenging choice sets and less diverse and lower variability in commuter profiles, carefully curated few-shot examples are responsible for nearly 80% of observed gains in the predictive performance, highlighting the importance of data context when model performance is initially weak. On the other hand, the joint contribution of Prompt-Type and Temperature remains below 2% in all datasets, indicating that these elements, while not irrelevant, are secondary considerations. Practitioners may use standard templates and moderate temperature settings without significant performance trade-offs. Therefore, prompt engineering and temperature tuning can be treated as last mile optimizations with marginal returns.
@@
With the model selection emerging as the dominant source of variation in predictive performance across datasets (Figure 4), it warrants a closer examination of individual model behaviour and their relative performance. Specifically, the model’s rank based on its weighted F1 scores under three learning styles across the different travel survey datasets. Figure 5 presents the aggregated ranks for each model, averaged across prompt styles and temperature settings for each configuration. In the zero-shot, where no contextual examples are provided to guide inference, DeepSeek R1 Distill Llama 8B and DeepSeek R1 Distill Qwen 7B consistently occupy the top ranks. Both models are distilled versions of the larger DeepSeek R1 reasoning model, employing reinforcement learning during distillation to preserve chain-of-thought capabilities in compact architectures of seven to eight billion parameters. Their performance illustrates how well designed distillation processes can embed advanced reasoning into lightweight, deployable models, making them highly suitable for tasks with no annotated demonstrations, and where minimum contextual guiding examples are available. Close competitors include Qwen 2.5 7B and Mistral 7B. These dense models incorporate advanced architectural innovations, such as rotary position encoding, SwiGLU activation functions, and grouped query or sliding-window attention mechanisms. Their ability to generalise under zero-shot settings points to the benefits of robust instruction tuning and efficient context handling in their pretraining.
@@
When random few-shot prompting is introduced, Gemma 3 12B emerges as the top performer. This large scale transformer leverages a hybrid attention mechanism that alternates between local and global attention blocks, allowing it to process long sequences up to 128,000 tokens without loss of coherence. Even with uncurated examples, the model demonstrates superior contextual absorption, significantly outperforming others. Nonetheless, the DeepSeek distillates and Qwen 2.5 7B remain competitive, reflecting the advantage of combining strong reasoning priors with scalable context length.
@@
In the targeted few-shot, where training examples are carefully selected based on similarity to the test case, Gemma 3 12B further consolidates its lead. Its architecture proves particularly effective at internalising structured demonstrations when available. Notably, Stealth 1.2 7B and Stealth 1.3 7B also rise in rank under this setting. These models, although less documented, are optimised for local inference and appear to have benefited from finetuning on diverse instructional datasets, allowing them to respond well to high quality contextual cues.
@@
This ranking analysis reinforces the importance of selecting models not solely by parameter count, but by architectural features and training history. The distilled DeepSeek variants offer high accuracy at low computational cost in zero-shot settings. Mid-sized dense models such as Qwen 2.5 7B and Mistral 7B provide a reliable trade-off between generalisability and efficiency, while larger models like Gemma 3 12B yield significant returns when prompted with curated examples. Overall, the results indicate that the optimal model choice depends critically on the deployed learning style and the availability of demonstration examples.

### v2

@@ 3.1.1 F1 Scores and Model’s Consistency
Figure 3 presents the distribution of weighted F1 scores achieved by each model across three travel survey datasets; Swissmetro, Brightwater, and London PMC under three learning styles: zero-shot, random few-shot, and targeted few-shot. The violin plots show each model’s distribution across four experimental runs, highlighting central tendency (mean), maximum performance (peak), and dispersion (IQR). The vertical span of each violin reflects the full value range across runs. This structure allows us to simultaneously evaluate performance levels and consistency across models and prompting styles while being able to assess the model’s sensitivity to varying prompting styles and temperature. For the remainder of this section, the results are presented and discussed dataset-wise, while Table 2 presents the summary, mean, peak and sensitivity/ consistency for top performing models.
@@
For the Swissmetro dataset, zero-shot experiments, DeepSeek R1 Distill Llama 8B outperforms others with a mean of 0.542 and peak of 0.568, maintaining a narrow IQR of 0.018, which indicates robust and insensitive performance under varying prompt conditions and temperature. Gemma 3 12B follows with a mean value of 0.497 and an even lower IQR of 0.011. Meanwhile, smaller models like Llama 3.2 1B record lower means (0.395) but also low dispersion (IQR = 0.011), which suggests that their output is rigid with limited prediction capacity at the base level, especially without any guiding examples.
@@
When transitioning to random few-shot learning, most models show gains in mean performance, though often at the cost of increased IQR or dispersion of performance results. This could mainly be attributed to the variability in the presented training profiles. DeepSeek R1 Distill Llama 8B reaches a new mean of 0.574 (peak = 0.617, IQR = 0.031), while Gemma 3 12B rises to 0.529 (peak = 0.558, IQR = 0.023), still maintaining less sensitivity to the variation of the presented training profiles, albeit higher than its baseline IQR performance.
@@
Targeted few-shot prompts significantly elevate both accuracy and stability. Gemma 3 12B reaches the highest overall mean of 0.619 (peak = 0.625), while tightening its IQR to just 0.0035, making it the most consistent top performer. DeepSeek Llama follows with a mean = 0.601 but with a higher sensitivity to the prompt style and temperature.
As for Brightwater, this travel survey dataset exhibits generally lower performance across all learning regimes. The complexity stems from its dynamic choice structure. Although each individual sample presents only 4 alternatives, the global choice set comprises 10 distinct modes. This variability, combined with relatively homogeneous commuter profiles, renders the prediction task more challenging. In the zero-shot setting, Gemma 3 4B achieves the highest mean score (0.345) and peak performance (0.393), but comparatively has a wide IQR of 0.080, indicating moderate sensitivity to prompt and temperature variations. DeepSeek R1 Distilled models, on the other hand, deliver a lower mean (0.310) but with higher consistency across the different runs (Average IQR = 0.020),
With random few-shot learning, incremental gains are observed across the board. Gemma 3 12B attains the highest mean of 0.405 and peak of 0.412 while achieving a tight IQR of 0.010. Llama 3.2 1B shows smaller improvements (mean = 0.311, peak = 0.321) with consistent spread (IQR = 0.008 ). Under the targeted few-shot, gains are noticed again with Llama 3 8B emerging as the top-performing model, achieving the highest mean (0.462) and peak (0.484) scores.
Models performance for London PMC shows a relatively higher baseline. In the zero-shot setup, Mistral 7B leads with a mean of 0.576 and a peak of 0.584, with IQR = 0.006. Stealth 1.3 7B performs less consistently, showing a much broader IQR = 0.171 despite competitive peaks. Random few-shot prompts elevate Gemma 3 12B to a mean of 0.661 (peak = 0.675), with an IQR of only 0.007. Meanwhile, Stealth models appear competitive with relatively high peak values, although with much higher sensitivity to the examples and/ or the variation in the prompting styles and temperatures. In the targeted few-shot, Gemma 3 12B acheives the highest mean (0.639) and a peak (0.647), while Qwen 2.5 7B hits the highest peak of 0.663 (mean = 0.622, IQR = 0.018). Stealth 1.3 7B here become more consistent (IQR = 0.012), and achieves a respectable overall predictive performance, which shows that they are more adaptable when prompted with tailored examples.
@@
Table 2: Top model performance summary per learning style across datasets
@@
Swissmetro
Zero-Shot 0.542 (DeepSeek R1) 0.568 (DeepSeek R1) 0.011 (Gemma 3 12B)
Random Few-Shot 0.574 (DeepSeek R1) 0.617 (DeepSeek R1) 0.023 (Gemma 3 12B)
Targeted Few-Shot 0.619 (Gemma 3 12B) 0.625 (Gemma 3 12B) 0.0035 (Gemma 3 12B)
Brightwater
Zero-Shot 0.345 (Gemma 3 4B) 0.393 (Gemma 3 4B) 0.008 (DeepSeek R1)
Random Few-Shot 0.405 (Gemma 3 12B) 0.412 (Gemma 3 12B) 0.010 (Gemma 3 12B)
Targeted Few-Shot 0.462 (Llama 3 8B) 0.484 (Llama 3 8B) 0.003 (Gemma 3 4B)
London PMC
Zero-Shot 0.576 (Mistral 7B) 0.584 (Mistral 7B) 0.006 (Mistral 7B)
Random Few-Shot 0.661 (Gemma 3 12B) 0.675 (Gemma 3 12B) 0.007 (Gemma 3 12B)
Targeted Few-Shot 0.639 (Gemma 3 12B) 0.663 (Qwen 2.5 7B) 0.012 (Stealth 1.3 7B)
@@
Across all three datasets, the transition from zero-shot to random few-shot prompting consistently yields notable improvements in mean weighted F1 scores, typically ranging from +0.03 to +0.07. However, this performance gain is frequently accompanied by an increase in interquartile range, with several models exhibiting expanded variability between runs. The pattern is mainly led by the sensitivity to the quality and coherence of uncurated in-context examples, which potentially undermines the consistency despite improvements in central tendency. Targeted few-shot prompting on the other hand, enhances predictive accuracy and systematically improves reliability. Selecting contextually relevant exemplars based on structured multi-layered similarity, IQR values are frequently reduced, often by 0.03 to 0.10, which highlights their stabilizing effect across models.
3.1.2 Variance Decomposition of Predictive Performance
@@
To further investigate the variability observed in the weighted F1 scores and identify which experimental factors most strongly influence predictive performance, we decompose the weighted F1 variance across the four key dimensions: the choice of language model (Model), the learning strategy (Shot-Type: zero-shot vs. few-shot), the prompt style (Prompt-style), and the model temperature (Temp). Figure 4 presents the decomposition of the variance share that can be explained by each factor across the three datasets. For each dataset, we fit an ordinary least squares (OLS) model of the form:
$\text{F1}_{\text{Weighted}}\sim C(\text{Model})+C(\text{Shot-Type})+C(\text{Prompt-Style})+C(\text{Temp})$ (30)
@@
We then perform a Type II ANOVA to assess the relative contribution of each factor. The sum of squares attributable to each term was normalized such that the four contributions sum to 100% of the explained variance.
@@
In Swissmetro, the choice of model is the dominant driver of performance variability, accounting for 70.6% of the explained variance. This indicates that the choice of a more capable model yields the greatest returns. The learning strategy or shot-type contributes 29.0%, which constitutes a significant portion of the predictive power. Prompting style and temperatures impact compared to model choice and learning strategy are marginal, explaining just 0.36% and 0.10% respectively. The analysis of the Brightwater dataset, on the other hand, exhibits a different pattern, with shot-type explaining a substantial 79.8% of the variance. This reflects the increased difficulty of the task and the impact of in-context examples in elevating performance. Here, model choice plays a secondary role (19.9%), while prompt-type and temperature again contribute minimally (0.21% and 0.12%). These findings suggest that for low-baseline tasks, few-shot examples can be transformative, whereas model choice or upgrades alone yield a less effective boost in returns.
As for the London PMC dataset, the analysis aligns more closely with Swissmetro. Model choice dominates (75.8% of variance), followed by Shot-Type (23.0%). Prompt-Type contributes slightly more here than in the other datasets (1.1%), though still marginal. Temperature influence is effectively null (0.005%). The results simply demonstrate the importance of selecting a performant LLM and a suitable learning strategy, while minor prompt adjustments or temperature tuning have limited practical value, and can be used for minor gains in terms of achieving optimal performance.
@@
Overall, the analysis shows that for more traditional choice sets and diversified commuter profiles through larger samples, upgrading the model explains over 70% of F1 variation as we can note in both Swissmetro and London PMC. However, for more challenging choice sets and less diverse and lower variability in commuter profiles, carefully curated few-shot examples are responsible for nearly 80% of observed gains in the predictive performance, hence the importance of data context when model performance is initially weak. On the other hand, the joint contribution of prompt-type and temperature remains below 2% in all datasets, indicating that these elements, while not irrelevant, are secondary considerations. Practitioners may use standard templates and moderate temperature settings without significant performance trade-offs. Therefore, prompt engineering and temperature tuning can be treated as last mile optimizations with marginal returns.
@@ 3.1.3 LLM’s Structure and Relative Performance Evaluation
With the model selection emerging as the dominant source of variation in predictive performance across datasets (Figure 4), it warrants a closer examination of individual model behaviour and their relative performance. Specifically, we rank the models based on their achieved weighted F1 scores under the three learning styles across the different travel survey datasets. Figure 5 presents the aggregated ranks for each model, averaged across prompt styles and temperature settings for each configuration. For zero-shot settings, where no contextual examples are provided to guide inference, DeepSeek R1 Distill Llama 8B and DeepSeek R1 Distill Qwen 7B consistently occupy the top ranks. Both models are distilled versions of the larger DeepSeek R1 reasoning model, employing reinforcement learning during distillation to preserve chain-of-thought capabilities in compact architectures of seven to eight billion parameters. Their performance proves how well designed distillation processes can embed advanced reasoning into lightweight, deployable models, making them highly suitable for tasks with no annotated demonstrations, and where minimum contextual guiding examples are available.
@@
When random few-shot prompting is introduced, Gemma 3 12B emerges as the top performer. This large scale transformer leverages a hybrid attention mechanism that alternates between local and global attention blocks, allowing it to process long sequences up to 128,000 tokens without loss of coherence. Even with uncurated examples, the model shows superior contextual absorption, significantly outperforming others. Nonetheless, the DeepSeek distillates and Qwen 2.5 7B remain competitive, mainly due to the advantage of combining strong reasoning priors with scalable context length.
@@
In the targeted few-shot, where training examples are carefully selected based on similarity to the test case, Gemma 3 12B further consolidates its lead. Its architecture proves particularly effective at internalizing structured demonstrations when available. Notably, Stealth 1.2 7B and Stealth 1.3 7B also rise in rank under this setting. These models, although less documented, are optimized for local inference and appear to have benefited from finetuning on diverse instructional datasets, allowing them to respond well to high quality contextual cues.
@@
This ranking analysis helps showcase the importance of selecting models not solely by parameter count, but by architectural features and training history. The distilled DeepSeek variants offer high accuracy at low computational cost in zero-shot settings. Mid-sized dense models such as Qwen 2.5 7B and Mistral 7B provide a reliable trade-off between generalizability and efficiency, while larger models like Gemma 3 12B yield significant returns when prompted with curated examples. Overall, the results indicate that the optimal model choice depends critically on the deployed learning style and the availability of demonstration examples.
@@
3.1.4 Learning Style and Learning Potential of LLMs for Predictive Tasks